OpenStudy (anonymous):

please help :( i will give medals MULTIPLE CHOICE: The measure of the largest angle of a certain triangle is 3 times the measure of the smallest angle. The measure of the remaining angle of the triangle is the average of the measures of the largest and smallest angles. Write and solve the system of equations that describes the measure of each angle of the triangle. a. 45 degrees, 60 degrees, 45 degrees b. 30 degrees, 60 degrees, 90 degrees c. 45 degrees, 45 degrees, 90 degrees d. 60 degrees, 60 degrees, 60 degrees.

OpenStudy (anonymous):

x + 3x +2x = 180 6x = 180 x = 30 3x = 90 (30+90)/2 = 60 b

OpenStudy (anonymous):

oh wow haha, i got it right before you sent me that @DSS thank you so much though!!

OpenStudy (anonymous):

Great!